OpenStudy (anonymous):

If sin(xy)=x^2, then dy/dx =

OpenStudy (anonymous):

implicit differentiation for sin(xy) differentiate and apply chain rule \[{d\over dx}\sin(xy)=\cos(xy){d\over dx}(xy)\]

OpenStudy (anonymous):

i did that and got \[\cos(xy)\times x \frac{ dy }{ dx } + y= 2x\]

OpenStudy (anonymous):

that should be the thing on the left

OpenStudy (anonymous):

\[{d\over dx}(xy)=x{dy\over dx}+y\] this result multiplies with the cosine thingy

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

ok.. \[ \cos(xy){d\over dx}(xy)=2x\\\text{but,} {d\over dx}(xy)=x{dy\over dx}+y\\\text{so,} \cos(xy)\left[x{dy\over dx}+y\right]=2x\\ x{dy\over dx}+y=2x\sec(xy)\\ x{dy\over dx}=2x\sec(xy)-y\\ {dy\over dx}=2\sec(xy)-{y\over x} \]

OpenStudy (anonymous):

oh!!! got it i